# The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals

### v1

@@ 5 Discussion
Privacy Expectations and Willingness to Donate.Privacy Expectations and Willingness to DonatePrivacy Expectations and Willingness to Donate. Privacy expectations significantly influence people’s willingness to donate health data: individuals are more willing to donate when their privacy expectations are met or exceeded, as presented in Figure 6. Thus, we argue that entities wishing to receive people’s data should not only focus on employing PETs but also on effectively communicating these measures to potential data donors. However, there is a risk of "privacy washing" [67, 26], where information about PETs and “vague claims” are used to raise expectations and mislead individuals into donating their personal data.
@@
The Role of Auditing.The Role of AuditingThe Role of Auditing. Auditing had a limited effect on raising privacy expectations, except in the case of purpose use restrictions in the for-profit scenario. Participants recognize that PETs are not flawless and that auditing does not practically mitigate the risk of data breaches. This result highlights a disconnect between the expert view, which often sees auditing as a fundamental guarantee to offer, and the participant view, which is more concerned with the overall protection of data rather than specific PETs implementations (e.g., a respondent writes “they can guarantee privacy all they want, things still get hacked”).
@@
Recommendations for Researchers.Recommendations for ResearchersRecommendations for Researchers. If researchers aim to foster trust in entities that collect personal health data, they may want to focus on effectively communicating about PETs that protect against data breaches. Technologies like multi-party computation, although not yet scalable, could enhance trust in data protection: they specifically distribute sensitive information across different entities to ensure that no single entity can reconstruct the information on its own, thereby minimizing the impact of a breach. However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved. We believe that convincing the public of the effectiveness of these measures presents an even greater challenge, as there is existing doubt and any failure of these systems could severely damage trust.
[…]
Recommendations for Regulators.Recommendations for RegulatorsRecommendations for Regulators. Our findings suggest the necessity of policies that ensure privacy practices align with public expectations. Given the strong assumptions people have about privacy protections, especially for non-profit entities, regulations should be designed to ensure that the protections assumed are consistently applied.

### v2

@@ 5. Concluding Discussion
Future Work.Future WorkFuture Work. In this work, we explicitly examined the influence of four privacy guarantees and two auditing guarantees. Future research could broaden this scope by incorporating a more diverse and comprehensive range of PETs, with a particular focus on understanding how multiple guarantees interact in real-world scenarios where they often coexist. Additionally, researchers can also conduct targeted studies to understand the perspectives of diverse demographic groups, including individuals with a strong distrust of data-handling practices. Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
Our study focused exclusively on individuals’ perspectives regarding formal medical data donation scenarios. However, medical data are also frequently collected through smart devices and donated with a single click via in-app prompts. Prior work emphasizes the role of modality and timing in the design space of privacy information (Schaub et al., 2015). Future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
Finally, prior work (Seberger and Patil, 2021) studying COVID-19 contract tracing apps – many created through a partnership between for-profits and non-profits – found that people assumed these apps would “operate according to the economics of surveillance capitalism…[making it] unreasonable to expect them to trust the system that profits from the sale of their data.” Future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
Appendix A Survey
In this section, we report the full text of our questionnaire. We first ask for a participant themselve or their close relatives have been diagnosed with a disease. In the rest of the questionnaire, we will choose at random one disease the participant has selected, if any, or one disease the participant has not selected. This is labeled in Italic as disease. The entity, also in Italic, may be either non-profit entity or for-profit entity, chosen at random.
Egocentricity
  1. (1)
    Have you ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
  2. (2)
    Have your partner and/or close relatives (parents, children, siblings) ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
Survey Introduction
Next, we present you with a scenario and ask you how you would respond. Please do your best to imagine yourself in this scenario. There is no right or wrong answer to these questions: we are interested to hear how you would feel about being in this scenario and what you would do.
Imagine that an entity entity wants to develop a new treatment for disease. They need medical data from people with and without disease to develop the treatment. They ask you to donate your medical record to help develop the treatment.
Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
Questions
We will now ask you a series of questions about this scenario. For your reference, the scenario text will appear at the top of each page.
  1. (1)
    How would you rate your understanding of the above scenario?
    • Fully understand
    • Mostly understand
    • Partially understand
    • Not understand
  2. (2)
    In this scenario how likely would you be to donate your medical record?
    • Very likely
    • Likely
    • Unlikely
    • Very unlikely
  3. (3)
    Why your are willing (or unwilling) to share your medical record with this entity?
  4. (4)
    Please rate your agreement with the following statement:
    I trust the entity will handle my data as described.
    • Strongly agree
    • Somewhat agree
    • Somewhat disagree
    • Strongly disagree
  5. (5)
    Please explain why you do (or do not) trust that the entity will handle your data as described.
  6. (6)
    In this scenario, how likely do you think it is that each of the following will occur?
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • The donated medical record will be deleted at a set point in time.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • The donated medical records will be used for another purpose without my consent.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • If I donate my data, I will meet Albert Einstein.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
  7. (7)
    History In real life, have you ever donated your medical record?
    • Yes, I have donated my medical record for disease
    • Yes, I have donated my medical record for research of other illnesses.
    • Yes, I have donated my medical record for non-medical purposes.
    • No, I was asked to donate my medical record but I declined to participate.
    • No, I have never been asked to donate my medical record.
    • I prefer not to answer.
Internet Skills
  1. (1)
    How familiar are you with the following terminology? Please choose a number between 1 and 5 where 1 represents “no understanding” and 5 represents “full understanding” of the term.
    • Advanced search
    • PDF
    • Spyware
    • Wiki
    • Cache
    • Phishing
  2. (2)
    Which of the following best describes your educational background or job field?
    • I have an education in, or work in, the field of computer science, computer engineering or IT
    • I DO NOT have an education in, nor do I work in, the field of computer science, computer engineering or IT.
    • Prefer not to answer.
  3. (3)
    Have you ever heard any of the following techniques? (select all that apply)
    • Differential Privacy
    • End-to-end encryption
    • Public-key cryptography
    • Trusted execution environment (TEE)
    • Pseudo-party computation
    • None of the above
Demographics
  1. (1)
    What year were you born in? (4 digits)
  2. (2)
    To which gender do you most identify? (Select all that may apply)
    • Woman
    • Man
    • Non-binary
    • Prefer not to answer
    • Prefer to self-describe
  3. (3)
    What is the highest level of education you have completed?
    • Less than high school degree (1)
    • High school graduate (high school diploma or equivalent including GED) (2)
    • Some college but no degree (3)
    • Associate’s degree (4)
    • Bachelor’s degree (5)
    • Advanced degree (e.g., Master’s, doctorate) (6)
    • Prefer not to answer (7)
Appendix B Code Book
@@
Data Donation.Data DonationData Donation.
  1. (1)
    Irrelevant response
  2. (2)
    Opt to donate regardless
  3. (3)
    Opt not to donate regardless
  4. (4)
    Require compensation
  5. (5)
    Only donate to a non-profit entity
  6. (6)
    No donation to a for-profit entity
  7. (7)
    Need more information
  8. (8)
    The provided protection is persuasive
  9. (9)
    Generally, the donation scenario looks good
  10. (10)
    Not satisfied with the provided protection
@@
Trustworthy.TrustworthyTrustworthy.
  1. (1)
    Trust without a specific reason
  2. (2)
    Irrelevant response
  3. (3)
    Perfect protection does not exist
  4. (4)
    The recipient entity has no reason not to protect the data
  5. (5)
    The recipient entity has no reason to protect the data
  6. (6)
    A non-profit entity is trustworthy
  7. (7)
    A for-profit entity is not trustworthy
  8. (8)
    Need more information
  9. (9)
    The provided protection is persuasive
  10. (10)
    Generally, the donation scenario looks good
  11. (11)
    Not satisfied with the provided statement